1.During normal beating, the heart creates a maximum 5.71 mV potential across 0.335 m of a person’s chest, creating a 1.01 Hz electromagnetic wave. What is the maximum magnetic field strength in the electromagnetic wave? Give your answer in picotesla (10-12T).
Answer #1

Electric field is given by

E=V/d =(5.71*10-3)/0.335 =0.01704 V/m

Maximum Magnetic field is given by

B=E/c =0.01704/(3*108)=56.8*10-12 T

B=56.8 pT